OpenStudy (anonymous):

Are the ratios 25/45 and 15/27 proportional? explain. I was trying to cross multiply can you please let me know what it is! Or at least how to! Thanks

Directrix (directrix):

If a/b = c/d, then a*d = b*c If these are proportional, 25/45 and 15/27, then 25/45 15/27, then 25*27 has to equal 45*15 Crank out the multiplications and see if they are proportional. @Toxic_Vapes